CFEC’s cash flow per share (CFPS) is currently $1.80 per share. Their 5-year average Price-to-Cash Flow per Share ratio is 16. If we expect their cash flow per share to grow by 8% next year, what would you expect their price to be next year?
Answer #1

CASH FLOW PER SHARE AFTER A YEAR= CFPS0*(1+g)^n

CASH FLOW PER SHARE AFTER 5 YEARS = 1.80*(1+0.08)^1 = 1.944

SO PRICE AFTER A YEAR = PRICE TO CASH FLOW PER SHARE X CASH FLOW PER SHARE AFTER A YEAR

PRICE AFTER A YEAR = 16 X 1.944 = 31.10 (rounded to 2 decimals)

Answer : 31.10 (Thumbs up please)